miller juice, Inc just paid a $3 dividend. The company is expected to pay a $3.50 dividend nest year and a $4 dividend in two years. After that, dividends are expected to grow 5% forever. If investors require a return if 12% on the investment what should Miller juice stock sell for today?
Answer #1
we have to use dividend discount model to compute the price of stock today
Price today = Present value of the future cash flow = PV of dividend + PV of terminal value
i ii iii iv=ii+iii v vi=v*iv
Year Dividend Terminal value Total cash flow PVIF @ 12% Present value
1 3.5 3.5 0.892857          3.13
2 4 60 64 0.797194       51.02
      54.15
Terminal value = Dividend in year 3/(required rate - growth rate)
=4*105%/(12%-5%)
60
therefore price today =       54.15